# Entanglement, Coherence, and Recursive Linking in Dicke states: A Topological Perspective

### v1

@@ 1 Introduction
The study of multipartite entanglement remains one of the most challenging and rewarding frontiers in quantum information science [thapliyal1999multipartite, horodecki2009, aczel2002, bengtsson2016brief, ainley2024multipartite]. While bipartite systems are well-understood through the lens of Bell states [bell1964einstein], systems comprising three or more qubits exhibit a significantly richer structure of correlations that cannot be classified by a single measure [dur2000, amico2008]. Historically, the classification of these states has relied on their transformation properties under Stochastic Local Operations and Classical Communication (SLOCC) [dur2000]. For example, in tripartite systems, the Greenberger-Horne-Zeilinger (GHZ) state [greenberger1990] and the W state [dur2000three] represent two inequivalent classes of genuine tripartite entanglement [dur2000]. However, a parallel and intuitive framework has emerged that seeks to understand these quantum correlations through the lens of topology, specifically, through the mathematical framework of knot theory [kauffman2016, adams1994]. P. K. Aravind pioneered this approach by drawing analogies between the fragility of the GHZ state and the Borromean rings [see Fig. (1)], where the removal of a single component disentangles the entire system [aravind1997, bhattacharyya2025symmetric].
@@
In this work, we extend this topological analysis to the family of Dicke states, denoted as $|D_{n}^{(k)}\rangle$. Originally introduced by Robert H. Dicke in 1954 in the context of quantum optics to describe superradiance [dicke1954coherence, bartschi2019deterministic, mukherjee2020preparing], these states have since found profound relevance in quantum information and multiqubit entanglement studies [li2008, toth2007detection]. Dicke states are a class of multi-qubit (or $n$ qubit) quantum states with a fixed number of excitations, representing a combinatorial superposition of computational basis states. These states are symmetric under the permutation of any two qubits. They are simultaneous eigenstates of the total angular momentum operator $J^{2}$ and $J_{z}$, with all qubits treated as spin-$\frac{1}{2}$ particles. Unlike the asymmetric $|Star\rangle$ state [cao2020fragility] or the fragile $|GHZ\rangle$ state, Dicke states exhibit a high degree of permutation symmetry and robustness, a property that has been verified in various experimental realizations using photonic qubits [agrawal2006perfect, kiesel2007experimental, wieczorek2009experimental]. We propose that the entanglement structure of $|D_{n}^{(k)}\rangle$ maps onto an $n$-component Hopf link (see Fig. (2), which is three-component Hopf link], where every qubit is pairwise linked to every other qubit. To rigorously establish this analogy, we employ two quantum quantifiers viz. (a) the Schmidt Rank, which diagnoses the presence of bipartite entanglement after particle loss (i.e. when measurement is made upon the system) [plenio2007], and (b) Quantum Coherence (specifically the $l_{1}$-norm) [cao2020fragility, roy2023exploring], which we interpret as the fluidity that maintains the topological links. We show that the recursive mathematical structure of Dicke states under projective measurement mirrors the recursive preservation of Hopf links, providing a concrete operational bridge between quantum optics and topology. The paper is organized as follows.
@@ 2 Theoretical Preliminaries
In this section, we outline the fundamental concepts necessary for our analysis. We review the definition of Dicke states, the formalism of projective measurements [nielsen2010quantum], and the two key metrics used to quantify the resulting states viz. Schmidt rank and Quantum Coherence. Finally, we establish the relevant topological definitions [rolfsen1976].
@@
where $\sum P_{j}$ denotes the sum over all distinct permutations of the qubits. As for example, when $n=3$ and $k=1$, the Dicke state is $|D_{3}^{(1)}\rangle$ which is $\frac{|001\rangle+|010\rangle+|100\rangle}{\sqrt{3}}$ i.e. the $|W\rangle$ state [dur2000three]. It is easy to check that swapping any two qubits leaves the state unchanged and hence it is permutation symmetric. Another example of Dicke state is inverted $|W\rangle$ state that corresponds to $|D_{3}^{(2)}\rangle$ [the inverted $|W\rangle$ state ($|\overline{W}\rangle$) is essentially the bit-flipped version of the $|W\rangle$ state and it is written as $|\overline{W}\rangle=X\otimes X\otimes X|W\rangle$, where $X$ is Pauli-$X$ gate]. The fundamental property of these states is their permutation symmetry while no single qubit plays a privileged role, contrasting sharply with that of graph states like the $|Star\rangle$ state.
@@ 2.2 Projective Measurements
A projective measurement on a single qubit is described by the orthogonal projectors $P_{0}=|0\rangle\langle 0|$ and $P_{1}=|1\rangle\langle 1|$ [nielsen2010quantum]. For a multipartite state $|\psi\rangle$, the probability of obtaining outcome $m\in\{0,1\}$ on a specific qubit is given by the Born rule:
@@
In our topological analysis, this measurement process is the quantum analogue of cutting [mukherjee2020preparing] or removing a ring from a link. The properties of $|\psi^{(m)}\rangle$ determine the stability of the remaining topological structure [radhakrishnan2024entanglement].
@@ 2.3 Schmidt Rank as an Entanglement Measure
To determine if the remaining qubits stay linked after a measurement, we utilize the Schmidt decomposition [peres1995, nielsen2010quantum]. A pure bipartite state $|\phi\rangle_{AB}$ can be written as:
@@ 2.4 Quantum Coherence
While Schmidt rank indicates the existence of entanglement, quantum coherence shows the strength of the superposition that supports it. It is to be noted that, Quantum Coherence is basis-dependent, i.e., when coherence is measured in different bases, the quantified value is different [roy2023exploring]. In this work, we have quantified coherence in the computational basis. To quantify coherence, we employ the $l_{1}$-norm measure of coherence. While multiple coherence measures exist, the $l_{1}$-norm offers both theoretical relevance and computational tractability. Now, for a density matrix $\rho$, the measure $l_{1}$- norm is defined as the sum of the absolute values of the off-diagonal elements [roy2023exploring]. This measure was rigorously quantified within the resource theoretic framework by Baumgratz et al. [baumgratz2014quantifying]. Denoting $l_{1}$- norm of coherence by $C_{l_{1}}(\rho)$, for a density operator $\rho$ with respect to a given state the said norm is defined as,
@@ 2.5 Topological Links
We utilize standard definitions from knot theory [adams1994, rolfsen1976] to classify the quantum states:
@@
  • Borromean Rings: A set of three rings where no two are linked, but the three together are inseparable. Cutting one releases the other two (analogous to the GHZ state [aravind1997, mukherjee2020preparing]).
@@
  • $n$-Hopf Link: A generalization where every component is pairwise linked to every other component. Cutting one component leaves a robust (n1)(n-1)-component structure intact. We posit this as the topological equivalent of the Dicke state [li2008, mukherjee2020preparing].
@@ 2.6.1 Link Fluidity via Quantum Coherence
In classical knot theory, a physical link is often viewed as a static, localized connection between components [adams1994]. However, in the quantum regime of Dicke states, the link is not a localized bond but a delocalized correlation distributed across the entire Hilbert space. We define Link Fluidity ($\mathcal{F}$) as the capacity of the entangled structure to redistribute its correlations dynamically upon the loss of a subsystem. Mathematically, this is quantified by the $l_{1}$-norm of quantum coherence, expressed in the computational basis as
@@
To facilitate our topological analysis, we introduce a theoretical distinction between two regimes of multipartite entanglement, governed by the magnitude of their quantum coherence. We propose that the stability of a topological link under measurement is determined by the fluidity of its support in the Hilbert space [cao2020fragility]. From Eq.(7), it is clear that the link fluidity or fluidity of a multipartite quantum state is directly proportional to coherence of the state.
  • Rigid Entanglement in the low coherence regime: We characterize states with very little quantum overlap as possessing rigid entanglement. In this scenario, the connection between particles relies on just a few specific links. It is like a bridge supported by only a single pillar. It is a brittle structure where everything depends on that one support holding up. Because there are no extra supports or backups, the structure cannot handle any local damage. Consequently, if you remove just one part (via measurement), the entire connection breaks, and the system instantly falls apart [radhakrishnan2024entanglement].
  • Fluid Entanglement in the high coherence regime: On the other hand, we characterize states with high quantum overlap as possessing fluid entanglement. In this regime, the connection is spread out across a vast network of many different links. It behaves less like a stiff chain and more like a spiderweb or a fluid mesh. This setup has high redundancy, meaning there are thousands of alternative pathways connecting the system. We believe this flexibility allows the connection to survive if one part is lost. Just as water flows to fill a gap, the remaining links redistribute the load, keeping the system connected even after a part is removed [roy2025environment].
@@ 2.7 The Operational-Topological Framework
    Linking (The Geometry): This provides the qualitative model. It describes the global connectivity pattern of the qubits (e.g., $n$-Hopf link vs. Borromean rings) and defines the system’s structural class [kauffman2016].
@@
    Projective Measurement (The Surgery): This serves as the operational probe [nielsen2010quantum]. In this framework, the quantum collapse of a single qubit maps directly to the topological deletion of a single loop. Thus we have
@@
    Schmidt Rank (The Indicator): This serves as the existence condition for the residual link after the surgery is performed [plenio2007].
@@
    Quantum Coherence (The Stability Measure): This serves as the mechanism of Link Fluidity. It quantifies the delocalization of the link across the Hilbert space, explaining why the connection survives the surgery [roy2023exploring].
@@ 3.1 Initial State Properties
We consider a general Dicke state describing a system of $n$ qubits sharing $k$ excitations. As defined in the literature [agrawal2006perfect] and our preliminary notes, the state $|D_{n}^{(k)}\rangle$ is the symmetric superposition of all distinct permutations of $k$ excitations ($|1\rangle$) and nkn-k ground states ($|0\rangle$).
@@ 4.1 Measurement Results and State Evolution:
The projective measurement yields two possible outcomes, $|0\rangle$ or $|1\rangle$. The post-measurement states of the remaining n1n-1 qubits preserve the Dicke structure but with modified parameters [radhakrishnan2024entanglement]. The complete analytical profile, including the Schmidt rank of the bipartite cut (qubit 1 vs. rest) and the $l_{1}$-norm of quantum coherence, is summarized in Table 1.
@@ 4.2.2 General Case ($0<k<n$)
Since the states $|D_{n-1}^{(k)}\rangle$ and $|D_{n-1}^{(k-1)}\rangle$ are orthogonal (their total excitation counts are different), this form constitutes a valid Schmidt decomposition [peres1995].
@@ 4.3 Quantum Coherence Analysis
Crucially, in the entangled regime ($0<k<n$), both residual coherence values are strictly positive, indicating that the remaining system retains quantum superposition [roy2025environment].
@@ 5 Topological Analogy: The Recursive $n$-Hopf Link
The defining feature of the Dicke state $|D_{n}^{(k)}\rangle$ is its complete permutation symmetry where every qubit contributes identically to the global entanglement structure. Translating this into topology, we view each qubit as a topological loop (or ring). Unlike asymmetric states such as the Star configuration [cao2020fragility], where specific qubits play central or peripheral roles, the Dicke state corresponds to a system of $n$ identical loops, all mutually interlinked with equal strength. We identify this structure as an $n$-Hopf link. In a topological sense, the entire system can be pictured as $n$ loops where there is no privileged connection between any specific pair; instead, the linking is uniform across the manifold [kauffman2016].
@@ 5.1.1 The Separable Regime ($k=0$ and $k=n$): Zero Link Density
  • Topological Interpretation: This corresponds to $n$ completely disjoint and unlinked loops. Each loop exists independently, with no crossings or linkages between them [adams1994].
@@ 5.1.2 The Linked Regime ($0<k<n$): The $n$-Hopf Link Structure
  • Topological Interpretation: The loops are now interlinked. The resulting structure is an $n$-Hopf-Linked configuration, where every pair is symmetrically linked [li2008].
@@ 5.2 Quantum coherence as link fluidity
While the Schmidt rank indicates that the links persist, the quantum coherence ($C_{l_{1}}$) explains how they persist. We have already introduced the concept of Link Fluidity in the sec. $2.6.1$ and it have been expressed in the Eq.(16) as $C_{l_{1}}(|D_{n}^{(k)}\rangle)=\binom{n}{k}-1$. In a static topological link (like a steel chain), the connection is localized [rolfsen1976]. However, in the quantum Dicke state, the linking is mediated by the delocalization of the $k$ excitations across the $n$ qubits. The coherence measure $C_{l_{1}}$ quantifies this delocalization.
@@ 5.3 Topological Stability and Recursive Preservation
A remarkable property of Dicke states is their stability under local measurements. Topologically, a projective measurement corresponds to cutting or removing one loop from the link configuration. For the $|GHZ\rangle$ state (analogous to Borromean rings [aravind1997]), cutting one ring destroys the entire structure. However, for the Dicke state, this operation does not destroy the overall linkage. Instead, it results in a smaller but topologically similar connectivity pattern. This recursive preservation gives rise to a self-similar hierarchy:

### v2

@@ 1 Introduction
The study of multipartite entanglement remains one of the most challenging and rewarding frontiers in quantum information science [30, 17, 1, 10, 4]. While bipartite systems are well-understood through the lens of Bell states [9], systems comprising three or more qubits exhibit a significantly richer structure of correlations that cannot be classified by a single measure [15, 5]. Historically, the classification of these states has relied on their transformation properties under Stochastic Local Operations and Classical Communication (SLOCC) [15]. For example, in tripartite systems, the Greenberger-Horne-Zeilinger (GHZ) state [16] and the W state [14] represent two inequivalent classes of genuine tripartite entanglement [15]. However, a parallel and intuitive framework has emerged that seeks to understand these quantum correlations through the lens of topology, specifically, through the mathematical framework of knot theory [18, 2]. P. K. Aravind pioneered this approach by drawing analogies between the fragility of the GHZ state and the Borromean rings [see Fig. (1)], where the removal of a single component disentangles the entire system [6, 11].
@@
In this work, we extend this topological analysis to the family of Dicke states, denoted as $|D_{n}^{(k)}\rangle$. Originally introduced by Robert H. Dicke in 1954 in the context of quantum optics to describe superradiance [13, 7, 21], these states have since found profound relevance in quantum information and multiqubit entanglement studies [20, 31]. Dicke states are a class of multi-qubit (or $n$ qubit) quantum states with a fixed number of excitations, representing a combinatorial superposition of computational basis states. These states are symmetric under the permutation of any two qubits. They are simultaneous eigenstates of the total angular momentum operator $J^{2}$ and $J_{z}$, with all qubits treated as spin-$\frac{1}{2}$ particles. Unlike the asymmetric $|Star\rangle$ state [12] or the fragile $|GHZ\rangle$ state, Dicke states exhibit a high degree of permutation symmetry and robustness, a property that has been verified in various experimental realizations using photonic qubits [3, 19, 32]. We propose that the entanglement structure of $|D_{n}^{(k)}\rangle$ maps onto an $n$-component Hopf link (see Fig. (2), which is three-component Hopf link], where every qubit is pairwise linked to every other qubit. To rigorously establish this analogy, we employ two quantum quantifiers viz. (a) the Schmidt Rank, which diagnoses the presence of bipartite entanglement after particle loss (i.e. when measurement is made upon the system) [24], and (b) Quantum Coherence (specifically the $l_{1}$-norm) [12, 28], which we interpret as the fluidity that maintains the topological links. We show that the recursive mathematical structure of Dicke states under projective measurement mirrors the recursive preservation of Hopf links, providing a concrete operational bridge between quantum optics and topology. The paper is organized as follows.
@@ 2 Theoretical Preliminaries
In this section, we outline the fundamental concepts necessary for our analysis. We review the definition of Dicke states, the formalism of projective measurements [22], and the two key metrics used to quantify the resulting states viz. Schmidt rank and Quantum Coherence. Finally, we establish the relevant topological definitions [27].
@@
where $\sum P_{j}$ denotes the sum over all distinct permutations of the qubits. As for example, when $n=3$ and $k=1$, the Dicke state is $|D_{3}^{(1)}\rangle$ which is $\frac{|001\rangle+|010\rangle+|100\rangle}{\sqrt{3}}$ i.e. the $|W\rangle$ state [14]. It is easy to check that swapping any two qubits leaves the state unchanged and hence it is permutation symmetric. Another example of Dicke state is inverted $|W\rangle$ state that corresponds to $|D_{3}^{(2)}\rangle$ [the inverted $|W\rangle$ state ($|\overline{W}\rangle$) is essentially the bit-flipped version of the $|W\rangle$ state and it is written as $|\overline{W}\rangle=X\otimes X\otimes X|W\rangle$, where $X$ is Pauli-$X$ gate]. The fundamental property of these states is their permutation symmetry while no single qubit plays a privileged role, contrasting sharply with that of graph states like the $|Star\rangle$ state.
@@ 2.2 Projective Measurements
A projective measurement on a single qubit is described by the orthogonal projectors $P_{0}=|0\rangle\langle 0|$ and $P_{1}=|1\rangle\langle 1|$ [22]. For a multipartite state $|\psi\rangle$, the probability of obtaining outcome $m\in\{0,1\}$ on a specific qubit is given by the Born rule:
@@
In our topological analysis, this measurement process is the quantum analogue of cutting [21] or removing a ring from a link. The properties of $|\psi^{(m)}\rangle$ determine the stability of the remaining topological structure [26].
@@ 2.3 Schmidt Rank as an Entanglement Measure
To determine if the remaining qubits stay linked after a measurement, we utilize the Schmidt decomposition [23, 22]. A pure bipartite state $|\phi\rangle_{AB}$ can be written as:
@@ 2.4 Quantum Coherence
While Schmidt rank indicates the existence of entanglement, quantum coherence shows the strength of the superposition that supports it. It is to be noted that, Quantum Coherence is basis-dependent, i.e., when coherence is measured in different bases, the quantified value is different [28]. In this work, we have quantified coherence in the computational basis. To quantify coherence, we employ the $l_{1}$-norm measure of coherence. While multiple coherence measures exist, the $l_{1}$-norm offers both theoretical relevance and computational tractability. Now, for a density matrix $\rho$, the measure $l_{1}$- norm is defined as the sum of the absolute values of the off-diagonal elements [28]. This measure was rigorously quantified within the resource theoretic framework by Baumgratz et al. [8]. Denoting $l_{1}$- norm of coherence by $C_{l_{1}}(\rho)$, for a density operator $\rho$ with respect to a given state the said norm is defined as,
@@ 2.5 Topological Links
We utilize standard definitions from knot theory [2, 27] to classify the quantum states:
@@
  • Borromean Rings: A set of three rings where no two are linked, but the three together are inseparable. Cutting one releases the other two (analogous to the GHZ state [6, 21]).
@@
  • $n$-Hopf Link: A generalization where every component is pairwise linked to every other component. Cutting one component leaves a robust (n1)(n-1)-component structure intact. We posit this as the topological equivalent of the Dicke state [20, 21].
@@ 2.6.1 Link Fluidity via Quantum Coherence
In classical knot theory, a physical link is often viewed as a static, localized connection between components [2]. However, in the quantum regime of Dicke states, the link is not a localized bond but a delocalized correlation distributed across the entire Hilbert space. We define Link Fluidity ($\mathcal{F}$) as the capacity of the entangled structure to redistribute its correlations dynamically upon the loss of a subsystem. Mathematically, this is quantified by the $l_{1}$-norm of quantum coherence, expressed in the computational basis as
@@
To facilitate our topological analysis, we introduce a theoretical distinction between two regimes of multipartite entanglement, governed by the magnitude of their quantum coherence. We propose that the stability of a topological link under measurement is determined by the fluidity of its support in the Hilbert space [12]. From Eq.(7), it is clear that the link fluidity or fluidity of a multipartite quantum state is directly proportional to coherence of the state.
  • Rigid Entanglement in the low coherence regime: We characterize states with very little quantum overlap as possessing rigid entanglement. In this scenario, the connection between particles relies on just a few specific links. It is like a bridge supported by only a single pillar. It is a brittle structure where everything depends on that one support holding up. Because there are no extra supports or backups, the structure cannot handle any local damage. Consequently, if you remove just one part (via measurement), the entire connection breaks, and the system instantly falls apart [26].
  • Fluid Entanglement in the high coherence regime: On the other hand, we characterize states with high quantum overlap as possessing fluid entanglement. In this regime, the connection is spread out across a vast network of many different links. It behaves less like a stiff chain and more like a spiderweb or a fluid mesh. This setup has high redundancy, meaning there are thousands of alternative pathways connecting the system. We believe this flexibility allows the connection to survive if one part is lost. Just as water flows to fill a gap, the remaining links redistribute the load, keeping the system connected even after a part is removed [29].
@@ 2.7 The Operational-Topological Framework
    Linking (The Geometry): This provides the qualitative model. It describes the global connectivity pattern of the qubits (e.g., $n$-Hopf link vs. Borromean rings) and defines the system’s structural class [18].
@@
    Projective Measurement (The Surgery): This serves as the operational probe [22]. In this framework, the quantum collapse of a single qubit maps directly to the topological deletion of a single loop. Thus we have
@@
    Schmidt Rank (The Indicator): This serves as the existence condition for the residual link after the surgery is performed [24].
@@
    Quantum Coherence (The Stability Measure): This serves as the mechanism of Link Fluidity. It quantifies the delocalization of the link across the Hilbert space, explaining why the connection survives the surgery [28].
@@ 3.1 Initial State Properties
We consider a general Dicke state describing a system of $n$ qubits sharing $k$ excitations. As defined in the literature [3, 25, 7] and our preliminary notes, the state $|D_{n}^{(k)}\rangle$ is the symmetric superposition of all distinct permutations of $k$ excitations ($|1\rangle$) and nkn-k ground states ($|0\rangle$).
@@ 4.1 Measurement Results and State Evolution:
The projective measurement yields two possible outcomes, $|0\rangle$ or $|1\rangle$. The post-measurement states of the remaining n1n-1 qubits preserve the Dicke structure but with modified parameters [26]. The complete analytical profile, including the Schmidt rank of the bipartite cut (qubit 1 vs. rest) and the $l_{1}$-norm of quantum coherence, is summarized in Table 1.
@@ 4.2.2 General Case ($0<k<n$)
Since the states $|D_{n-1}^{(k)}\rangle$ and $|D_{n-1}^{(k-1)}\rangle$ are orthogonal (their total excitation counts are different), this form constitutes a valid Schmidt decomposition [23].
@@ 4.3 Quantum Coherence Analysis
Crucially, in the entangled regime ($0<k<n$), both residual coherence values are strictly positive, indicating that the remaining system retains quantum superposition [29].
@@ 5 Topological Analogy: The Recursive $n$-Hopf Link
The defining feature of the Dicke state $|D_{n}^{(k)}\rangle$ is its complete permutation symmetry where every qubit contributes identically to the global entanglement structure. Translating this into topology, we view each qubit as a topological loop (or ring). Unlike asymmetric states such as the Star configuration [12], where specific qubits play central or peripheral roles, the Dicke state corresponds to a system of $n$ identical loops, all mutually interlinked with equal strength. We identify this structure as an $n$-Hopf link. In a topological sense, the entire system can be pictured as $n$ loops where there is no privileged connection between any specific pair; instead, the linking is uniform across the manifold [18].
@@ 5.1.1 The Separable Regime ($k=0$ and $k=n$): Zero Link Density
  • Topological Interpretation: This corresponds to $n$ completely disjoint and unlinked loops. Each loop exists independently, with no crossings or linkages between them [2].
@@ 5.1.2 The Linked Regime ($0<k<n$): The $n$-Hopf Link Structure
  • Topological Interpretation: The loops are now interlinked. The resulting structure is an $n$-Hopf-Linked configuration, where every pair is symmetrically linked [20].
@@ 5.2 Quantum coherence as link fluidity
While the Schmidt rank indicates that the links persist, the quantum coherence ($C_{l_{1}}$) explains how they persist. We have already introduced the concept of Link Fluidity in the sec. $2.6.1$ and it have been expressed in the Eq.(16) as $C_{l_{1}}(|D_{n}^{(k)}\rangle)=\binom{n}{k}-1$. In a static topological link (like a steel chain), the connection is localized [27]. However, in the quantum Dicke state, the linking is mediated by the delocalization of the $k$ excitations across the $n$ qubits. The coherence measure $C_{l_{1}}$ quantifies this delocalization.
@@ 5.3 Topological Stability and Recursive Preservation
A remarkable property of Dicke states is their stability under local measurements. Topologically, a projective measurement corresponds to cutting or removing one loop from the link configuration. For the $|GHZ\rangle$ state (analogous to Borromean rings [6]), cutting one ring destroys the entire structure. However, for the Dicke state, this operation does not destroy the overall linkage. Instead, it results in a smaller but topologically similar connectivity pattern. This recursive preservation gives rise to a self-similar hierarchy:
@@ 6 Conclusion
7 Conflict of Interest:
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
8 Data Availability Statement:
Our manuscript has no associated data.